# AutoRAN: Automated and Zero-Touch Open RAN Systems

### v2

@@ VI-B LLM-Based Deployment
In this section, we evaluate our LLM-based deployment pipeline. To do so, we use Claude 3.7 Sonnet to generate a series of 33 different deployment prompts. Each prompt provides some specific requirements in NL and is associated with the required element(s) in the deployed network. For example, if the query is “generate a 5G network based on GPU acceleration”, the associated requirement is that the DU-low uses ARC. We evaluate the deployment pipeline by running each prompt 10 times, for different sizes of LLMs. The LLM is deployed on one of the control-plane nodes, which are equipped with a NVIDIA L40S GPU with $40$ GB of VRAM. Figure 8(a) shows that the larger model is significantly better at providing satisfactory configurations, with Qwen2.5:32B reaching up to $80$% correct responses, and with correctness consistently decreasing as we decrease the size of the model. Furthermore, Figure 8(b) shows that this extra performance does not necessarily come at the cost of a larger runtime, with Qwen2.5:1.5B having the largest runtime despite its low success-rate. We observe (see Figure 8(c)) that this concurs with the high number of iterations in the initial phase (where we verify if values are missing, see Figure 7): the model struggles to find any valid configuration until it times out. On the other hand, larger models have similar iteration counts due to their ability to output a correct configuration (in the sense that we almost always obtain a deployable network, but that network is not necessarily the one the user asked for, as per Figure 8(c)).
@@ VI-C End-to-End Deployment and Testing Workflows
In this section, we show experimental results to validate the automated end-to-end test workflows described in Section V, where we use Sierra Wireless 5G modems as UEs.
@@
Together, the following five bars show the average time required to perform a complete test with five different gNB combinations: (i) GH200 server with ARC; (ii) Gigabyte E251 server with ARC; (iii) Gigabyte E251 server without ARC; (iv) Microway server with OAI and USRP; and (v) Microway server with srsRAN and USRP. The Test Specification bar shows the time taken from when the user specifies which deployment and test configuration to use, to when the gNB Flask API is called to start the test. The pull time, included in this value, is dependent on the size of the image and the network speed. Having a local and private registry inside OpenShift allows AutoRAN to pull the image at the maximum speed allowed by the connection between the cluster and the NAS, that is up to $10$ Gbps. Extra time is still required to the various layers of the image in the target node. On the other hand, after the image has been pulled at least once on a node, it is thereafter kept in its cache registry, and subsequent pulls are in the order of a few milliseconds. Therefore, by pre-loading the image in the internal cache on the node, the pull time becomes negligible even for the bigger images, such as cuBB (around $40$ GB).
After the image has been deployed in a pod on the targeted node, the gNB software needs to be instantiated (gNB setup in Figure 10). As shown, the time to run the gNB on the GH200 and Gigabyte servers is quite similar, at around $18$ s. This time accounts for the time to initialize libraries, GPU and NIC for the L1 acceleration. On the other hand, since to run OAI without ARC on the Gigabyte server only one container is needed and the gNB is monolithic, the start time is shorter, about $8$ s. As the RUs is always in idle state (i.e., in the on state but without actively transmitting or receiving) and initialized outside of OpenShift with via M-plane-like functionalities before the gNB pod deployment (see Section IV), we do not include initialization (approximately $60$ s for a Foxconn RU) and reconfiguration times in the instantiation time of CU/DU. This is different from the USRPs deployment, where initialization times are negligible. Indeed, the time taken to start the gNB and connect it to the USRP for both srsRAN and OAI is around $10$ s. Finally, the last two bars show the time taken for the UE (Sierra Wireless 5G modem in our case) to connect to the gNB, and the time taken to collect the results and send them to the data collector pod respectively. It is worth mentioning that we do not report the duration of the data transmissions between gNB and UE, as it depends on the value specified in the test specifications (e.g., run an iPerf test for $60$ s, see Listing 5).
@@ VI-D1 Performance of the Automated Testing Workflow
Figures 12 and 13 show the downlink and uplink TCP throughput achieved during the specific end-to-end (E2E) experiments performed through the workflows described in Section V to test five different combinations of gNBs with a single Sierra Wireless as the UE. For this specific test, we select TCP as a way to measure not only raw performance but also the realistic goodput reaching the considered UE.
@@ VI-D2 Performance Comparison of NVIDIA ARC
Our results show that the added virtualization and integration into AutoRAN does not lead to relevant performance difference in raw throughput when deploying NVIDIA ARC gNBs. It is worth mentioning that the cuBB logs show some delayed slots from time to time in the OpenShift-based implementation (probably due to the missing optimized kernel), but this situation does not seem to affect performance. The AutoRAN approach shows therefore how it is becomes possible to apply infrastructure sharing to the RAN world as in the cloud computing one while almost the same performance
VI-E RAN Coexistence with Generic Workloads
In this section, we test whether generic workloads on a RAN oriented node has any effect on the RAN itself. We want to evaluate this coexistence capabilities for various reasons, including building of containers on the targeted node and any other workloads that the OpenShift scheduler allocates on that node itself on the shared CPUs.
We show how running increased CPU-based consumption for a long period of time does not affect the throughput of the L1. We use the RuSIM emulation in Figure 15. We simulate high demanding workloads using the Linux tool stress on either $10,20,30$ isolated CPUs and on the generally available shared CPUs of the system.
@@ Acknowledgments
The authors would like to express their gratitude to Jing Xu, Anupa Kelkar, and Chris Dick from NVIDIA Corporation for their feedback on NVIDIA ARC.

### v3

@@ VI-B LLM-Based Deployment
In this section, we evaluate our LLM-based deployment pipeline. To do so, we use Claude 3.7 Sonnet to generate a series of 33 different deployment prompts. Each prompt provides some specific requirements in NL and is associated with the required element(s) in the deployed network. For example, if the prompt is “generate a 5G network based on GPU acceleration”, the associated requirement is that the DU-low uses ARC. We evaluate the deployment pipeline by running each prompt 10 times, for different sizes of LLMs. The LLM is deployed on one of the control-plane nodes, which are equipped with a NVIDIA L40S GPU with $40$ GB of VRAM. Figure 9(a) shows that the larger model is significantly better at providing satisfactory configurations, with Qwen2.5:32B reaching up to 8080% correct responses, and with correctness consistently decreasing as we decrease the size of the model. Furthermore, Figure 9(b) shows that this extra performance does not necessarily come at the cost of a larger execution time (measured from the time the user sends the prompt to the time the input to the Tekton pipeline is generated, see Fig. 7), with Qwen2.5:1.5B having the largest runtime despite its low success-rate. We observe (see Figure 9(c)) that this concurs with the high number of iterations in the initial phase (where we verify if values are missing, see Figure 7): the model struggles to find any valid configuration until it times out. On the other hand, larger models have similar iteration counts due to their ability to output a correct configuration (in the sense that we almost always obtain a deployable network, but that network is not necessarily the one the user asked for, as per Figure 9(c)).
@@ VI-C End-to-End Deployment and Testing Workflows
In this section, we show experimental results to validate the automated E2E test workflows described in Section V, where we use Sierra Wireless 5G modems as UEs. We leave the extension of the automated tests to smartphone UEs for future work, while we use these devices for manual and mobile tests.
@@
Together, the following five bars show the average time required to perform a complete test with five different gNB combinations: (i) GH200 server with ARC; (ii) Gigabyte E251 server with ARC; (iii) Gigabyte E251 server without ARC; (iv) Microway server with OAI and USRP; and (v) Microway server with srsRAN and USRP. The Test Specification bar shows the time taken from when the user specifies which deployment and test configuration to use, to when the gNB Flask API is called to start the test. The pull time, included in this value, is dependent on the size of the image and the network speed. Having a local and private registry inside OpenShift allows AutoRAN to pull the image at the maximum speed allowed by the connection between the cluster and the NAS, that is up to $10$ Gbps. Extra time is still required to the various layers of the image in the target node. On the other hand, after the image has been pulled at least once on a node, it is thereafter kept in its cache registry, and subsequent pulls are in the order of a few milliseconds. Therefore, by pre-loading the image in the internal cache on the node, the pull time becomes negligible even for the bigger images, such as cuBB (around $40$ GB). After the image has been deployed in a pod on the targeted node, the gNB software needs to be instantiated (gNB setup in Figure 10). As shown, the time to run the gNB on the GH200 and Gigabyte servers is quite similar, at around $18$ s. This time accounts for the time to initialize libraries, GPU and NIC for the L1 acceleration. On the other hand, since to run OAI without ARC on the Gigabyte server only one container is needed and the gNB is monolithic, the start time is shorter, about $8$ s. As the RUs is always in idle state (i.e., in the on state but without actively transmitting or receiving) and initialized outside of OpenShift with via M-plane-like functionalities before the gNB pod deployment (see Section IV), we do not include initialization (approximately $60$ s for a Foxconn RU) and reconfiguration times in the instantiation time of CU/DU. This is different from the USRPs deployment, where initialization times are negligible. Indeed, the time taken to start the gNB and connect it to the USRP for both srsRAN and OAI is around $10$ s. Finally, the last two bars show the time taken for the UE (Sierra Wireless 5G modem in our case) to connect to the gNB, and the time taken to collect the results and send them to the data collector pod respectively. It is worth mentioning that we do not report the duration of the data transmissions between gNB and UE, as it depends on the value in the test specifications (e.g., run an iPerf test for $60$ s, see Listing 5). Finally, multiple tests can be executed either as independent tests (e.g., one for iPerf with TCP and one for iPerf3 in UDP) where the gNB is restarted prior to the execution of the next test, or consecutively without redeploying the pod.
@@ VI-D1 Performance of the Automated Testing Workflow
Figures 12 and 13 show the downlink and uplink TCP throughput achieved during the specific E2E experiments performed through the workflows described in Section V to test five different combinations of gNBs with a single Sierra Wireless as the UE. For this specific test, we select TCP as a way to measure not only raw performance but also the realistic goodput reaching the considered UE.
@@ VI-D2 Performance Comparison of NVIDIA ARC
Our results show that the added virtualization and integration into AutoRAN does not lead to relevant performance difference in raw throughput when deploying NVIDIA ARC gNBs. It is worth mentioning that the cuBB logs show some delayed slots from time to time in the OpenShift-based implementation (probably due to the missing optimized kernel), but this situation does not seem to affect performance. Therefore, the AutoRAN approach shows how it is possible to apply infrastructure sharing to the RAN world as in the cloud computing one without significant performance degradations. As at the time of this writing, open-source RAN stacks are not optimized for Ultra Reliable and Low Latency Communications (URLLC) use cases, we evaluate latency by exchanging ICMP packets between a Samsung Galaxy S23 Commercial Off-the-Shelf (COTS) UE and core network on AutoRAN as well as on the bare-metal NVIDIA ARC deployment. Results for this evaluation are shown in Figure 15. We notice that the performance between the original NVIDIA ARC and the AutoRAN deployments are similar, with small variations in latency related to non-deterministic channel conditions and an average ICMP Round Trip Time (RTT) of $18$ ms.
Figure 15: CDF plot of achieved ICMP latency of a COTS UE.
VI-E Coexistence with Different Workloads
In this section, we test whether generic workloads on a RAN oriented node has any effect on the RAN itself. We want to evaluate this coexistence capabilities for various reasons, including building of containers on the targeted node and any other workloads that the OpenShift scheduler allocates on that node itself on the shared CPUs. We show how running increased CPU-based consumption for a long period of time does not affect the throughput of the L1. We use the RuSIM emulation in Figure 16. We simulate high demanding workloads using the Linux tool stress on either $10,20,30$ isolated CPUs and on the generally available shared CPUs of the system.
@@ VI-E Coexistence with Different Workloads
VI-F AutoRAN Resilience
The OpenShift container orchestration platform AutoRAN is based upon—which is, in turn, based on Kubernetes—natively provides high-availability mechanisms to compensate for node failures by “evicting” (i.e., moving) workloads from unresponsive to functioning cluster nodes. As mentioned in Section III, however, RAN workloads typically require specific hardware functionalities to be available on the compute nodes. Therefore, they can only move across nodes where these are available (in practice, across nodes part of the same MCP), while an AutoRAN-orchestrated redeployment is needed across nodes with different MCPs. As shown in Fig. 17, the eviction happens automatically after a timeout that starts when OpenShift detects the node to be unavailable. The duration of the timeout is configurable between 0 s (the pod gets evicted as soon as the node appears unhealthy) and infinite (the pod never gets evicted), with a default of 5 minutes. Further analysis of this parameter is left for future work. Further, due to limitations of the protocol stack implementations we leverage, a new instantiation of the pipeline is required to re-start the gNB once the workload is re-allocated. This triggers an RU reset to accommodate the new DU MAC address (alternatively, the DU MAC can be spoofed, but this does not represent a scalable solution without additional middleware, as discussed in [37]). Note that, during this process, UEs can detect a Radio Link Failure (RLF) and connect to a different cell, or directly reconnect to the new gNB, depending on the duration of the eviction timer.
Figure 17: Workflow of AutoRAN when a node is detected to be unavailable.
@@ Acknowledgments
The authors would like to express their gratitude to Jing Xu, Anupa Kelkar, and Chris Dick from NVIDIA Corporation for their feedback on NVIDIA ARC. The authors disclose that AI tools (ChatGPT, Claude) have been used to perform minor edits and reformulations of the text.